### v4

# Log-normal Superstatistics in the Confined Motion of Ants

Abstract
We report the emergence of Log-normal Superstatistics in the collective motion of ants confined in a quasi-2D arena and exposed to a panic-inducing stimulus. A data-driven superstatistical Langevin model accurately reproduces the transition from stationary behavior to an organized escape response, characterized by non-Gaussian velocity distributions and a fluctuating diffusion coefficient. Our findings show that danger information propagates via a memory-limited, cascade-like mechanism, resulting in a stable cluster formation despite individual memory constraints. These discoveries establish a crucial connection between Superstatistics formalisms and living active matter beyond a unicellular level, and provide a foundation for the understanding of the biological origin of Log-normal type diffusion in confined environments.
Complex systems often exhibit emergent behavior that challenges classical Statistical Mechanics. In particular, far-from-equilibrium systems can feature fast local dynamics that coexist with slow fluctuations of intensive parameters such as temperature or diffusivity [1, 2, 3, 4, 5]. Superstatistics (i.e., a superposition of different local equilibrium statistics) has been successfully applied to describe these frameworks with strong parameter fluctuations, such as turbulence, convection and active particles [6, 7, 8, 9]. In living matter, devoted studies include DNA architecture and the motion of proteins and tumor cells [10, 11, 12]. Although its relevance to macroscopic biological systems has been theoretically proposed [13], it had not been directly explored experimentally. The collective behavior of animals, particularly eusocial insects like ants, presents an ideal substrate for applying superstatistical frameworks. Colonies of ants respond to environmental stimuli using distributed sensing and local communication, displaying pattern formation, foraging trails, and escape dynamics [14, 15, 16, 17].
A canonical starting point for modeling directional animal movement is the Langevin equation [18, 19, 20, 21, 13]:
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma\xi_{i}(t)$ (1)
where $v_{i}$ is the $i$-th Cartesian component of velocity of each individual animal, $\gamma$ is a positive parameter that encloses directional persistence, and $\xi_{i}$ is a Gaussian white noise of intensity $\sigma$. Superstatistics extends this view by allowing intensive parameters such as $\beta=2\gamma/\sigma^{2}$ to fluctuate on longer time scales. The system is assumed to reach local equilibrium over a short time $\tau_{1}$, while the parameter $\beta$ varies over a slower scale $\tau_{2}\gg\tau_{1}$. Consequently, within an interval of duration $\tau_{2}$, $\beta$ can be treated as approximately constant and the velocity distribution takes the Gaussian form $p(v_{i}|\beta)=(\beta/2\pi)^{1/2}\exp{(-\beta v_{i}^{2}/2)}$ . Therefore, to capture the long-time behavior, one integrates over the ensemble of local equilibria weighted by the distribution $f(\beta)$, yielding the marginal density for the $i$-th velocity component $f(v_{i})=\int_{0}^{\infty}p(v_{i}|\beta)f(\beta)\,d\beta$ which yields diverse non-Gaussian stationary distributions documented in many physical systems [22, 23, 24, 25, 26, 27].
Here, we ask whether a fundamental statistical framework underlies ants movement in confinement and persists or adapts when ants are perturbed by panic stimuli. We show that such a framework aligns to the principles of Superstatistics and study the motion of confined ants as a case study for superstatistical analysis. Ants display non-Gaussian velocity distributions and fluctuating diffusivity, features seen in single-particle tracking of soft and biological matter [28, 29, 30, 31, 32, 33, 34]. When exposed to danger signals transmitted from a neighboring arena, ants collectively form a cluster that shifts away from the source of danger, revealing a transition from short individual memory to long-term collective behavior. We propose a super-statistical model that captures both regimes, providing what appears to be a unique experimental report of Superstatistics in the animal kingdom, and, notably, the Log-normal type in living systems.
Experiments were performed in dual acrylic arenas, labelled danger and safe, of $6\times 6\times 1\,\mathrm{cm^{3}}$ connected via a Thin barrier designed to allow contacts of different types (see End Matter) (Fig. 1a). Each cell was populated with nearly 20 adult workers of the Cuban ant species Atta insularis (Fig. 6a), randomly selected from a single colony and previously studied in both natural and laboratory settings [35, 36, 37, 38, 39]. After a 5-minute baseline, $50\,\mathrm{\mu L}$ of insect repellent was introduced into the danger cell. Video data were recorded for a total of 15 minutes at 30 fps. In control (Ctrl) experiments no ants were introduced into the danger arena. Each protocol was repeated 10 times using different individuals to avoid memory effects [40].
A significant shift of the center of mass (CM), measured via simple video binarization, away from the barrier was observed after introducing the repellent, indicating an escape response (Fig. 1b,d). The negligible shift in the Ctrl protocol (Fig. 1c,e) confirmed that this effect was not due to passive diffusion but driven by inter-ant communication. The CM velocity increases sharply after repellent exposure, indicating a rapid initial reaction followed by a slower phase in which ants form a stable cluster away from the barrier. This two-stage process, i.e., early activation followed by clustering, will be discussed later. Additional protocols, time-resolved dynamics, statistical validation, and experiment videos with CM evolution are provided in End Matter and Supplementary Information (SI), where we also indicate that optical signals or simple antennation are not significant mechanisms for panic transmission (Fig. 7).
We now show that single-ant velocities and accelerations can be analyzed and incorporated into a Langevin-based model that accurately reproduces the ensemble’s CM dynamics in the safe cell before repellent injection in Ctrl experiments, where ants interact only with each other and the barrier, free from influence by ants in the danger arena (baseline regime). This unperturbed state establishes a reference model of ant behavior under confinement based on the Langevin equation (1), which we later extend to incorporate interactions with ants in the danger cell and the transmission of danger information in Thin experiments (panic regime) (Fig. 1d).
Figure 1: (a) Experimental setup and snapshots of experiments for (b) Thin and (c) Ctrl protocols with the center of mass (CM) trajectory colored by 5-minute intervals (green, blue, red). (d) Net horizontal CM displacement relative to its initial position for the Thin protocol, as predicted by the superstatistical model in Eq. (5) (SSModel); (e) corresponding experimental data for the Ctrl protocol.
Baseline regime was characterized by tracking 18 randomly selected ants, using the yupi library for trajectory analysis [41]. To avoid boundary effects, only trajectory segments beyond $\ell_{w}=0.6\,\mathrm{cm}$ from all walls were used in the analysis, where $\ell_{w}$ is a characteristic length obtained from the exponential decay of ant acceleration near walls (Fig. 5).
[…]
Fig. 2a shows that the acceleration component parallel to the velocity averages to $\mu(v)=-\gamma v$, where $\gamma^{-1}=0.53\,\mathrm{s}$. The orthogonal component averages to zero across speeds. However, the intensity of fluctuations in both components depends nonlinearly on speed and follows the same trend within their error bars. We model these behavior with
$\sigma(v)=\Sigma+\Delta\Sigma\,(1-e^{\displaystyle{-v/v_{\sigma}}})$ (2)
where $\Sigma$ captures resting fluctuation intensity, $\Delta\Sigma$ sets the total increase in noise intensity from rest to high speeds, and $v_{\sigma}=1.99\,\mathrm{cm/s}$ modulates the speed dependence. Although alternative forms (e.g., power laws) were tested, the exponential fit performed best, avoiding unphysical values like $\sigma(0)\leq 0$. Furthermore, the temporal correlation of ant velocities decays exponentially (Fig. 8b), yielding a persistence time $\gamma^{-1}$ consistent with the slope of the parallel acceleration.
Figure 2: (a) Mean (blue/orange) and fluctuation intensity (dark/light green) of parallel and perpendicular acceleration components vs. speed. Parallel acceleration scales linearly, perpendicular remains near zero, and both speed-dependent standard deviations fit Eq. (2). Note that $\mu(v)$ and $\sigma(v)$ differ in units (i.e., $\mathrm{cm\,s}^{-2}$ and $\mathrm{cm\,s}^{-3/2}$, respectively). (b) Autocorrelation function of acceleration fluctuations around their mean (Noise Autocorrelation Function, NACF). Parallel and perpendicular components are shown in blue and orange, respectively. The fluctuations exhibit no temporal correlation, indicating uncorrelated noise. (c) Probability density function (PDF) of parallel and perpendicular acceleration fluctuations. The non-Gaussian shape is accurately captured by the SSModel (Eq. 4), which combines a Log-normal and Gaussian dynamics, but it cannot be explained only by Gaussian noise (G. Noise) with the state-dependent intensity $\sigma(v)$. (d) Time-average of local flatness (Eq. (3)). The dashed line marks the Gaussian value. Inset: typical realization of $D(t)$ (dark red) varies more slowly than $v_{x}(t)$ (dark blue) over $\sim 50\,\mathrm{s}$.
The acceleration fluctuations show negligible temporal correlations (Fig. 2b). However, Fig. 2c shows that they deviate from the typical Gaussian noise of Eq. (1), and that although state-dependent noise $\sigma(v)\xi_{i}(t)$ yields non-gaussianity, it fails to explain our observations. Thus a natural step beyond constant noise intensity is to treat it as a slowly varying stochastic process, placing us in the realm of Superstatistics. Therefore, these non-Gaussian fluctuations can be explained based on a superposition of local Gaussians with variances $\beta(t)$ fluctuating over long temporal scales. In consequence, to verify time scale separation in our data, we computed the time-dependent flatness $\kappa_{i}(t)$ of velocity components [43]:
$\kappa_{i}(t)=\frac{1}{T-t}\int_{0}^{T-t}\frac{\langle(v_{i}-\langle v_{i}\rangle)^{4}\rangle}{\langle(v_{i}-\langle v_{i}\rangle)^{2}\rangle^{2}}dt^{\prime}$ (3)
Here, $T$ is the total measurement time, and $\langle\cdot\rangle$ denotes a local mean over window $t$. The slow time scale $\tau_{2}$ was extracted as the window size at which the flatness $\kappa_{i}(t)$ reaches its Gaussian value, indicating local equilibrium of velocity statistics (Fig. 2d). We find $\tau_{2}=3.3\,\mathrm{s}$, which is much longer than the velocity decorrelation time $\tau_{1}=\gamma^{-1}=0.53\,\mathrm{s}$; their ratio $\tau_{2}/\tau_{1}\approx 6$ confirms the existence of an intensive parameter $\beta(t)$ that fluctuates on a timescale about 6 times slower than the velocity itself.
Figure 3: (a) Probability density (PDF) of diffusion coefficient $D$ (standardized) fitted to a Log-normal distribution. Inset shows the tail up to 15 standard deviations. (b) Velocity PDF with Gaussian, Laplace, and power-law (Lévy stable) fits. The SSModel (4) (solid line) captures the central peak and tails beyond Gaussian decay.
The slowly varying stochastic process $\beta(t)$ was then computed over sliding windows of length $\tau_{2}$ as the inverse of the local velocity variance, $\beta(t)=[\langle v_{i}^{2}(t)\rangle_{\tau_{2}}-\langle v_{i}(t)\rangle_{\tau_{2}}^{2}]^{-1}$. Consequently, we explored suitable powers $p$ such that $\beta^{p}$ could make the dynamics fall into one of the universality superstatistical classes (e.g., $\chi^{2}$ or Log-normal [44]) and found that p=1p=-1 yields the diffusion coefficient $D\propto\beta^{-1}$ as the relevant slowly varying quantity. Fig. 3a shows that $D$ follows a Log-normal distribution with shape parameter $s=1.27$. Although Log-normal and stretched exponential fits can appear similar when tails are limited [44], the inset shows a robust Log-normal fit extending over 15 standard deviations. Typical time series of $v_{x}(t)$ and $D(t)$ in Fig. 2d (inset) visually confirm their separate time scales.
Once gathered the measured persistence time $\gamma^{-1}$, the speed-dependent noise intensity $\sigma(v)$, and the slowly varying diffusion coefficient $D(t)$, we set them into a superstatistical Langevin model [2] for the velocity component $v_{i}(t)$ in the safe cell before panic onset:
$\frac{d}{dt}v_{i}(t)=-\gamma v_{i}(t)+\sigma(v)\sqrt{D(t)}\,\xi_{i}(t)$ (4)
We refer to this as the SSModel. For simulations, at each time step, $D(t)$ and the Gaussian noise $\xi_{i}(t)$ are combined into $\eta(t)=\sqrt{D(t)}\,\xi_{i}(t)$, which is standardized and scaled by $\sigma(v)$. Numerical integration was performed using a Runge-Kutta method of a strong order 1.5 within a Milstein scheme to compute Stratonovich-type integrals [45]. Fig. 3b shows that the SSModel accurately predicts the velocity distribution: no discrepancies are observed in the shapes, consistent with the arena’s spatial homogeneity and isotropy. The model reproduces both the low-speed peak and the non-Gaussian tails through the combined effects of $\sigma(v)$ and Log-normal fluctuations in $D(t)$. It also captures the non-Gaussian fluctuations of parallel and perpendicular accelerations (Fig. 2c).
Panic regime was examined by tracking 48 trajectories of ants during the next 10 minutes following repellent injection. Several observables remained unchanged (Fig. 8): mean acceleration stayed linear in speed, $\tau_{2}$ remained stable, and both $\gamma^{-1}$ and $D$ showed no significant changes. Centered acceleration distributions also retained their shape, indicating preserved local Gaussian equilibria for $t<\tau_{2}$. However, the velocity distributions developed stretched tails, consistent with gradual clustering in the safe cell, and non-gaussianity persisted due to the interplay of $v$-dependent noise and Superstatistics. The only clearly altered observable was $\sigma(v)$, which maintained isotropy but evolved over time. Specifically, two parameters, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$, underwent a sharp, brief increase at minute $9.2$, then relaxed, while $\Sigma(t)$ stayed constant (Fig. 4a). For later simulations, $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ were fit to contiguous back-to-back exponential functions with matching growth and relaxation rates, selected for their simplicity in characterizing rapid changes. The timing at which this peak occurs will subsequently be used as a key measure of one of the most significant timescales in the ants’ mechanism in response to panic.
As previously noted, the CM begins to shift almost immediately after repellent introduction (Fig. 1d), with fluctuations masking the earliest rise but a clear upward trend by minute 6, suggesting a response of about 1 minute. A biologically grounded mechanism for this rapid response is recruitment-based information transfer, as observed in tandem communication in ants [46], and it may proceed as follows. Initially dispersed, ants near the barrier likely receive danger cues quickly upon panic induction at minute 5, and become “activated”. Each then inform to another nestmate after a characteristic time $\tau_{f}$, doubling the number of activated ants in a cascade that ends at $T=1\,\mathrm{min}$ with $N_{T}=20$ informed individuals. This implies $\dot{N}=N/\tau_{f}$, with $N(0)=1$, yielding $\tau_{f}=T/\ln N_{T}=20\,\mathrm{s}$. While multiple ants may simultaneously access the barrier, leading to overlapping information flows, we simplify this dynamics as being driven by a single “effective ant” that triggers the cascade, i.e., a representation of the collective’s fast, memory-limited reaction to panic cues.
When all ants are informed at minute 6, a slower dynamics arises. It can be characterized by the time between the moment when all members were activated and the peak observed in Fig. 4a at minute $9.2$, i.e., by a characteristic time $\tau_{s}=3.2\,\mathrm{min}$. Hence, an extended SSModel based on Eq. (4) that incorporates boundary effects and panic dynamics mediated by the times $\tau_{f}$ and $\tau_{s}$, was evaluated as follows:
$\begin{split}\frac{d}{dt}v_{i}(t)&=-\gamma(v_{i}(t)-\tilde{v}_{i}(t))\\ &\quad\,+\sigma(v,t)\sqrt{D(t)}\,\xi_{i}(t)+F_{w}(r_{i})\end{split}$ (5a)
$\sigma(v,t)$ $=\Sigma(t)+\Delta\Sigma(t)\,(1-e^{\displaystyle{-v/v_{\sigma}(t)}})$ (5b)
$F_{w}(r_{i})$ $=a_{0}(e^{\displaystyle{-r_{i}/\ell_{w}}}-e^{\displaystyle{-(r_{i}-L)/\ell_{w}}})$ (5c)
$\tilde{\mathbf{v}}(t)$ $=\left(\frac{F_{w}(\tilde{x})}{\gamma}\frac{1-e^{\displaystyle{-(t-t_{p})/\tau_{s}}}}{1-e^{\displaystyle{-(t-t_{p})/\tau_{f}}}},\;0\right)$ (5d)
Here, $r_{i}$ represents the $x$-$y$ position of an ant, and $F_{w}(r_{i})$ in Eq. (5c) mimics wall forces with intensity $a_{0}=1.0\,\mathrm{cm/s^{2}}$, as fitted in Fig. 5a, and arena size $L=6.0\,\mathrm{cm}$. The drift velocity $\tilde{\mathbf{v}}(t)$ in Eq. (5d) has zero $y$ component due to the directionality of the panic effect, while its $x$ component reflects fast and slow dynamics with time scales $\tau_{f}$ and $\tau_{s}$. It is modeled as an asymmetric sigmoid centered at $t_{p}=5.5\,\mathrm{min}$ and saturates at $F_{w}(\tilde{x})/\gamma$, balancing wall repulsion and panic drive. The only free parameter, $\tilde{x}$, corresponds to the final cluster position. Fig. 1d shows this extended SSModel reproducing CM motion using experimentally derived parameters, with all values listed in Table 1.
Figure 4: (a) Time evolution of the parameters in Eq. (2). $\Delta\Sigma(t)$ and $v_{\sigma}(t)$ show a sharp change at minute 9.2, while $\Sigma(t)$ remains stable. (b) Probability of finding an ant from the safe cell at a distance less than or equal to $2\,\mathrm{cm}$ from the barrier. The last 5 minutes of the experiment are highlighted in orange, where the curve remains approximately constant near zero. During this interval, the ant cluster is stable.
The diffusion coefficient not only fluctuates on similar timescales before and after panic, but also retains its Log-normal statistics, suggesting a fundamental, stress-resistant feature of confined ant interactions, as reflected in stable acceleration PDFs (Fig. 8d). Likewise, the near-constancy of $\Sigma(t)$ indicates that danger does not alter the baseline noise intensity, while the return of $\Delta\Sigma(t)$ to its initial value implies that total noise intensity $\Sigma+\Delta\Sigma$ is effectively “remembered”. This superstatistical framework can be narrated in three stages of response to panic. First, from minute 5 to 6, a cascade begins as ants near the barrier become informed and transmit the danger signal across the safe cell, with a characteristic time $\tau_{f}=20\,\mathrm{s}$ comparable to the ant memory timescale reported in the literature [15]. In the second stage, from minute 6 to 9.2, the drift phase emerges as the CM shifts rightward as informed ants increasingly bias their movement away from the barrier, with a timescale $\tau_{s}=3.2\,\mathrm{min}$. Finally, from minute 9.2 to 15, the cluster stabilizes as the drift velocity $\tilde{\mathbf{v}}(t)$ saturates, reaching a force balance between the rightward panic repulsion and the leftward mechanical repulsion from the opposite wall.
A remaining issue is the contrast between the long-term stability of the cluster despite ants’ short individual memory. One possible explanation is that the cluster is not entirely static: some ants periodically leave the cluster, explore the barrier, and return (see video at SI). We hypothesize that the information feedback provided by these individuals is rapidly transmitted to the rest of the densely-packed ants, so the cluster-formation behavior is constantly reinforced. This idea is supported by Fig. 4b, which shows a drop but not a disappearance in barrier exploration probability following repellent injection. As the cluster stabilizes, its left edge remains roughly $2\,\mathrm{cm}$ from the barrier. The sustained but reduced traffic suggests that returning ants subtly refresh the danger signal, helping maintain cluster cohesion during the final 5 minutes of the experiment.
In summary, we have shown that Log-normal Superstatistics emerges as a robust statistical framework that captures the transition from normal to danger behavior in confined ants, while maintaining continuity and providing a foundation upon which panic-induced dynamics can be superimposed. We reported non-Gaussian behavior, a slowly varying diffusion coefficient, and a collective escape response triggered by inter-ant communication. A data-driven superstatistical model with Log-normal diffusion successfully reproduced these features and captured the danger transition, relying on a single free parameter. The model uncovers a fast, memory-limited information cascade followed by a slower collective reorganization into a stable cluster.
Acknowledgements from A. Reyes to the Abdus Salam ICTP STEP program for financial support. We thank A. Celani, D. Gordon, R. Mulet, T. Tohme and D. Machado for valuable discussions, G. Simon for the 3D-printed of the cells, and A. Serrano for assistance with Fig. 1. E. Altshuler acknowledges an Invited Professorship supported by CNRS at the University of Lyon.
Appendices
Figure 5: (a) Acceleration as a function of position (component wise). Inset: acceleration absolute value as a function of the distance from the walls. A characteristic length $\ell_{w}=0.6\,\mathrm{cm}$ is defined as the effective wall thickness. (b) The grey sections have been removed from a sample trajectory, as they are within a distance $\ell_{w}$ from the walls: in order to ensure that statistics is not contaminated by boundary effects, we only use for analysis the non-gray sections.
I. Boundary effects: The baseline regime was used to evaluate boundary effects while measuring the extent of wall “forces”. Experimental acceleration components (i.e., $a_{x}$ and $a_{y}$) were arranged based on increasing values of $x$- and $y$- position components, respectively. Both acceleration time series were binned and a running average was taken (main panel of Fig. 5a). The ants slow down as they approach the walls, revealing the influence of the boundaries as if they were confined by a soft, square-box-like potential. Acceleration absolute value against the distance from the wall revealed a characteristic length $\ell_{w}=0.6\,\mathrm{cm}$, obtained by fitting an exponential function (inset of Fig. 5a). Henceforth, analysis was conducted with trajectory segments that were consistently at least $\ell_{w}$ units away from every boundary, to discard boundary effects (Fig. 5b).
II. Mechanisms for danger transmission: To isolate potential mechanisms of danger transmission, we used three barrier types (Fig. 6b): a transparent Glass wall allowing only visual cues; a Thick barrier restricting the interaction to mutual antennation; and a Thin barrier allowing full contact between ants in adjacent arenas (Fig. 6c).
Figure 6: (a) Ant used in experiments: leafcutter ant Atta insularis, a species native to Cuba known as “bibijagua”. (b) Sketch of the three types of barriers used for the protocols Glass, Thick and Thin. The latter is also used for the Ctrl protocol. (c) Top view of ants near a thick and thin barriers. Short distance contacts are not allowed by the thick barrier. (d) Lateral view of an ant near a thin barrier. As a reference, the length of the ants bodies is around $1\,\mathrm{cm}$.
To characterize collective motion across protocols, we tracked the ants’ CM in both arenas (Fig. 7a-d). In all cases, the CM in the danger cell remained near the geometric center regardless of barrier type. The same held for the safe cell, except under the Thin protocol, where the CM systematically shifted away from the barrier after repellent was introduced (Fig. 7c). Fig. 7e-h shows the time evolution of the horizontal CM position in the safe cell, averaged over 10 experimental repetitions. A clear displacement emerges only in the Thin case after the injection at minute 5. In contrast, the Ctrl case (i.e., where the danger cell is empty) shows no comparable motion, indicating that passive diffusion alone cannot account for the observed shift.
Figure 7: Snapshots of experiments (a) Glass, (b) Thick, (c) Thin and (d) Ctrl. Green, blue, and red traces show the center of mass (CM) trajectory for consecutive 5-minute windows. (e-h) Horizontal CM evolution in the safe cell, relative to its initial position. Shading shows standard deviation. (i) Net horizontal displacement of the ants CM based on 10 repetitions of each type of experiment.
To quantify this, we computed the net CM displacement in the safe cell as the average difference in $x$-coordinate between minutes 5 and 15 of each experiment (Fig. 7i). We performed one-tailed Mann-Whitney U tests to evaluate the null hypothesis that each pair of samples (Thin vs. Glass, Thick, and Ctrl) originates from equal populations and used a significance level of $\alpha=0.05$. We obtained $U=84.0$, $89.0$, and $71.0$, and corresponding $p$-values of $0.0008$, $0.0002$, and $0.0187$, respectively. Since all $p$-values fall below the threshold, the null hypothesis is rejected in each case.
These results indicate that only the Thin barrier enables the transmission of panic-inducing information sufficient to trigger a significant escape response in ants within the safe cell. The Ctrl data reinforce this, showing that repellent diffusion alone, without ant-to-ant signaling, has no effect. Minimal shifts in the Glass and Thick protocols further suggest that neither visual cues nor simple antennation are effective mechanisms for danger transmission, but touching other body parts seems necessary.
Table 1: Values for model parameters fitted to data.
$\ell_{w}$ $a_{0}$ $\gamma^{-1}$ $\Sigma$ $\Delta\Sigma$ $v_{\sigma}$
$\mathrm{(cm)}$ $\mathrm{(cm/s^{2})}$ $\mathrm{(s)}$ $\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(cm\,s^{-3/2})}$ $\mathrm{(cm/s)}$
0.6 1.0 0.53 0.34 1.99 1.99
$s$ $\tau_{f}$ $\tau_{s}$ $L$ $t_{p}$ $\tilde{x}$
$\mathrm{(s)}$ $\mathrm{(min)}$ $\mathrm{(cm)}$ $\mathrm{(min)}$ $\mathrm{(cm)}$
1.27 20 3.2 6.0 5.5 3.8
Figure 8: (a) Speed dependence of parallel and perpendicular acceleration components across time intervals. (b) Local flatness vs lag time, computed over trajectory segments from time intervals $[0,5)$ to $[14,15)$ minutes. Final time of each interval is shown in legend. (c) Velocity autocorrelation functions (VACF) vs lag time. The solid line has slope $\gamma=(0.53\,\mathrm{s})^{-1}$. Notice that the slow time scale $\tau_{2}$ is not captured by the VACF. (d) Probability density function (PDF) of centered parallel and perpendicular components of acceleration fluctuation. (e) PDF of Cartesian velocity components. (f) Temporal and speed dependence of the intensity of acceleration fluctuations for the parallel (blue) and perpendicular (orange) components after panic was induced. Continuous lines show fits to Eq. (2).